Jane, Paul, and Peter can finish painting the fence in two hours. If Jane does the job alone she can finish it in 5 hours. If Paul does the job alone he can finish it in 9 hours. How long will it take for Peter to finish the job alone?
Answer by mananth(16949)


Jane, Paul, and Peter can finish painting the fence in two hours.
All the three can do 1/2 the job in 1 hour
Jane does the job alone can finish it in 5 hours.
she can do 1/5 of the job in 1 hour

Paul does the job alone can finish it in 9 hours.
He can do 1/9 of the job in 1 hour
Let Peter take x hours to do the job alone
So he does 1/x of the job in 1 hour
1/5 + 1/9+1/x = 1/2
1/x = 1/2 - (1/5 +1/9)
1/x = 1/2 - 14/45
1/x = (45-28)/90
1/x = 17/90
x = 90/17
Peter can do the job alone in 5.3 hours